# spacing=nonfrench

### v1

@@ Abstract
Generative Artificial Intelligence (GenAI) browser assistants integrate powerful capabilities of GenAI in web browsers to provide rich experiences such as question answering, content summarization, and agentic navigation. These assistants, available today as browser extensions, can not only track detailed browsing activity such as search and click data, but can also autonomously perform tasks such as filling forms, raising significant privacy concerns. It is crucial to understand the design and operation of GenAI browser extensions, including how they collect, store, process, and share user data. To this end, we study their ability to profile users and personalize their responses based on explicit or inferred demographic attributes and interests of users. We perform network traffic analysis and use a novel prompting framework to audit tracking, profiling, and personalization by the ten most popular GenAI browser assistant extensions.
We find that instead of relying on local in-browser models, GenAI browser assistants largely depend on server-side APIs, which can be invoked automatically without explicit user interaction. When invoked, these GenAI browser assistants collect and share webpage content, often the full HTML DOM and sometimes even the user’s form inputs, with their first-party servers. Some GenAI browser assistants also share identifiers and user prompts with third-party trackers such as Google Analytics. The collection and sharing continues even if a webpage contains sensitive information such as health or personal information such as name or social security number entered in a web form. We find that several GenAI browser assistants infer demographic attributes such as age, gender, income, and interests and use this profile—which carries across browsing contexts—to personalize responses. In summary, our work shows that GenAI browser assistants can and do collect personal and sensitive information for profiling and personalization with little to no safeguards.
@@ 1 Introduction
With rapid advancements in Natural Language Processing
over the past few years, Large Language Models (LLMs) have grown in scale to incorporate billions of parameters to significantly enhance their contextual understanding and language generation capabilities [29]. As the foundational technology behind many Generative Artificial Intelligence (GenAI) systems, LLMs have played a pivotal role in driving the growth of GenAI across diverse applications such as real-time user assistance, automated content generation, and interactive chatbots. Besides providing a more conversational and personalized search experience due to their search capabilities, they also support agentic capabilities. For instance, they can make a reservation on behalf of the user by performing appropriate searches and filling out the form.
Popular search engines are already integrating LLMs to enhance their search performance – for example, Google Search displays a Gemini-generated overview [42] in response to the user’s search query, while Microsoft Bing’s Copilot uses OpenAI’s model [28]. However, search engines are limited in scope to analyzing a user’s activity on their own platform, restricting their access into user’s broader browsing behavior across the web. To address this limitation, a class of assistants, which we refer to as “GenAI browser assistants” have emerged to leverage the capabilities of generative models in improving a user’s overall browsing on the Internet.
GenAI-based browser assistants are packaged as extensions that can be installed onto a user’s browser, providing them access to everything that a user does in their browser. More specifically, existence in context of an extension improves the ability of these assistants to track a user’s click, websites browsed, content accessed, queries asked, and personal information entered online by storing identifiers to associate user state across websites. The user’s online browsing activities act as a context to browser assistants to build a profile on them. As a result, although the assistants provide a highly personalized browsing experience, this affects user privacy. An online user may inadvertently end up sharing sensitive information regarding themselves, without realizing the resultant consequences. The shared information could be stored on remote servers of the browser assistants or shared with third-parties to target ads to the user.
Despite operational costs reaching as high as \$700,000 per day [22], most LLM platforms do not yet rely on ads for monetization. However, the industry is seeing a shift towards in-house advertising, as observed with Perplexity AI’s sponsored follow-up questions [46]. This raises user privacy concerns similar to those in the existing online advertising ecosystem. With these developments, our web is moving towards an era of in-browser Generative AI – a browser integrated with a local LLM model so as to maintain privacy by retaining the user’s data within the browser and yet guaranteeing the same performance quality as that of a cloud-based LLM model. Chrome’s Built-in AI is in its origin trial and would allow websites and web applications to perform AI-powered tasks within the user’s browser through APIs such as Prompt API [12], eliminating their need to deploy or manage its own AI models. GenAI-based browser assistants aim to provide a similar experience by either directly or indirectly relying on cloud-based LLM models to personalize user’s browsing experience. As a result, it is important to understand the privacy risks that arise from these GenAI-based browser assistants at the cost of user profiling and personalization. Millions of users already use these assistants to personalize their browsing experience on a daily basis. Despite such a large user base, it is unknown how exactly these browser assistants are designed, if, when, what, and how they handle user-dependent information to personalize their responses and the risks they pose to user privacy. To this end, we propose a novel framework to systematically audit GenAI-based browser assistants by addressing the following research questions:
@@
RQ1. How is the architecture of GenAI browser assistants designed? To this end, we qualitatively analyze network traffic activity across different extensions to reason the following design choices: (1) backend model capabilities, (2) response architecture, (3) context restrictions, and (4) response variability.
RQ2. Do GenAI browser assistants collect and share user information in response to their queries? In order to assess privacy risks, it is necessary to analyze how user data is handled. Users may either inadvertently expose personal information to the browser assistants while they browse private spaces on the web; or browser assistants may actively collect user’s browsing related data while they use them in their day-to-day lives. We focus on analyzing network traffic to study both sides of user tracking – (1) implicit and (2) explicit collection and sharing of user data with first-party or third-party entities.
RQ3. Are GenAI browser assistants capable of profiling a user based on their browsing behaviour to personalize its responses? We define profiling as the process of storing and analyzing the collected user data, including their preferences, behaviors, and characteristics, to build a holistic view about the user. While leveraging the profiled information to create an individualized experience is regarded as personalization. We propose a novel prompting framework that – first, leaks and tests for five user attributes – location, age, gender, wealth, and interests for profiling, and second, understands if GenAI browser assistants use leaked information to personalize their responses.
Our research provides the following key contributions with respect to each of the research questions described previously:
@@
  • Architecture. We analyze 10 of the most popular GenAI browser assistants and find that all-but-one uses server-side response generation, while only one operates client-side. We also observed some assistants to automatically invoke response generation based on the user’s search query. Furthermore, 8 assistants were observed to isolate contexts across page navigations, while the remaining two shared user prompt history across different navigation-based browsing contexts.
@@
  • User Tracking. Usage of assistants while browsing through private online spaces resulted in collection of full DOM to partial webpage by different assistants. Merlin was observed to extract even webform inputs, resulting in collection and sharing of medical health records from university health portal (hem.ucdavis.edu), student academic records (from canvas.edu), and SSN (entered on sa.www4.irs.gov/wmr/). Sider and Merlin shared chat and user identifiers with google-analytics.com while TinaMind shared them with analytics.google.com. Merlin also shared user’s raw query with Google analytics, suggesting potential for tracking and retargeting across Google platforms. User’s chat history from past conversations was also shared with first-party servers of four assistants. Moreover, Harpa and Copilot stored the full history in the background service worker’s IndexedDB storage, suggesting its persistence across browsing sessions.
@@
  • Personalization. For search experiments, four extensions (ChatGPT for Google, Copilot, Monica, and Sider) demonstrated profiling for all five user attributes regarding location, age, gender, wealth and interests as well as in-context and out-of-context profiling and personalization prompts. Two of them (Perplexity and TinaMind) didn’t show strong signals of profiling or personalization while Harpa only showed in-context personalization.
@@ 2 Background
In this section, we provide background on GenAI systems. In Section 2.1, we describe evolution, functions and challenges of these systems. Next, in Section 2.2, we discuss how GenAI browser assistants work and functionalities they offer.
2.1 Generative AI Systems
The evolution of GenAI. Artificial Intelligence has a history spanning over 50 years, dating back to the 1960s when rule-based systems were first introduced [15]. Over the years, rule-based models were improved using statistical machine learning approaches and multi-layered neural networks that allowed for development of natural language processing applications. Traditional language models relied on sequential processing that made it difficult to capture long-range dependencies in data. The advent of transformer architectures in 2017 [47] addressed this limitation by allowing all elements of the sequence to be processed simultaneously, enabling the model to capture relationships between words regardless of their position in the sequence. This innovation, coupled with massive datasets, improved computational power, and model architecture advancements led to the development of large language models (LLMs) like the generative pre-trained transformer (GPT) [40].
@@ 2.1 Generative AI Systems
How do GenAI systems work? Generative AI systems, like large language models, function on iteratively predicting the “next” token in a sequence, based on the context provided by previous tokens, to inform its predictions [54]. The underlying transformer-based architecture allows self-attention mechanism to weigh the importance of different words in a sequence differently [47]. When a user performs a query, the model encodes the input into an embedding space and generates a probability distribution over potential next tokens. It selects a token based on this distribution and repeats this process until the output is complete. Following this approach, GenAI systems are able to generate coherent and context-aware text.
@@
Challenges with GenAI systems. Despite their phenomenal generative capabilities, GenAI systems face several challenges that impact their performance, usability, and broader applicability. It is important to take these into account to truly understand how GenAI systems function. The first fundamental issue lies in its “probabilistic” nature. As aforementioned, the prediction of the next token relies on a probabilistic distribution, resulting in different outputs even for the same input, making these genAI systems less predictable than the deterministic ones. This variability is influenced by the “temperature” parameter that controls the randomness of generated output. The value of this parameter typically ranges between 0 and 1 – where higher values result in selection of less probable tokens, either producing more creative responses or outputs that are less coherent or relevant [38]. A lower temperature, on the other hand, results in more deterministic responses. The second challenge is the high cost of training these models. OpenAI’s GPT-4 cost up to \$78M, while Google’s Gemini cost upto \$191M – 15 times higher than the training cost of its precursor models [7]. This limits the frequency at which these models are trained. Moreover, model training leverages large-scale training data to improve its ability to generate more accurate and enhanced responses to user queries. However, it fails to incorporate recent information (e.g., recent news) that was not part of the training process without access to live search capabilities. Lastly, while general-purpose models excel at a wide-range of functions, adapting GenAI systems to domain-specific applications like personalized assistants demands significant effort as well as resources. These standalone models have access to just the user-performed queries on LLM platform’s website limiting their context to only what is explicitly provided by the user in their prompts when they visit their platform website. This makes personalization challenging.
@@ 2.2 Generative AI Browser Assistants
How do GenAI browser assistants function? To fully leverage capabilities of GenAI models to personalize user experience on the web, a category of GenAI systems – Gen AI browser assistants have recently emerged. These are essentially browser extensions that function as wrappers on the top of open-source models like OpenAI’s ChatGPT [36], Google’s Gemini [16], Meta’s Llama [26], etc. They record activities that a user performs on the web and queries the open-source model in the backend as per the user’s need to provide a more personalized response. Distinct from the foreground user activities, these browser assistants operate in a separate context as a service worker that is always running in the background when the extension is active. However, extension’s background code can inject Javascript snippets called ‘content scripts’ in every page that a user visits in the foreground. These content scripts can further request and include additional scripts to support necessary extension functionalities such as JQuery for simplified DOM manipulation for instance. Content scripts can also contact background script to share logged user activities. These capabilities provide GenAI browser assistants with kind of a super access to everything that a user does in their browser. For example, assistants can monitor all searches made by the user in their browser, access what content they seek daily on the web, etc. Monitoring user journey across the web, helps browser assistants suggest what a user might be looking for ‘in-the-moment’ by incorporating a highly specific context that is needed for better personalization. For instance, writing assistants can adapt to an individual’s writing style to auto-generate email replies. In summary, GenAI browser assistants address numerous challenges described above related to generic GenAI systems to truly personalize user’s online experience and enhance user interaction and productivity.
@@
Functionalities offered by GenAI Browser Assistants. The need for GenAI browser assistants stems from the growing complexity of digital ecosystems, where users are often overloaded with too much information and require personalized and efficient solutions to manage their needs as they navigate through the web. Based on the offered functionalities, we classify GenAI browser assistants into 3 broad categories – search-based assistants, content-based assistants, and automation assistants. Search-based assistants primarily improve user’s browsing by offering features such as natural language responses to search queries. They also support web page summarization, interactive Q&A with webpages, and text highlighting. Content-based assistants manage content-specific tasks such as writing tasks, social media post generation, meeting transcript summarization, YouTube video explanation, SEO generation, etc. Automation assistants focus on automating workflows to streamline tasks. They include tools that aid in auto-filling forms, coding and debugging agents, task scheduling, automated data extraction and scraping, voice-enabled assistance, etc. All these assistants thus act as intelligent intermediaries to provide a wide array of functionalities to enhance user experience through context-aware suggestions, making them more intuitive, efficient, and tailored to individual needs.
3 Related Works
@@
In this section, we provide literature overview of related works across two dimensions – privacy issues in extensions and privacy issues in GenAI models. We conclude by discussing how our work compares to the prior research in this space.
@@
Privacy Issues in Extensions. Despite enhanced user experience, browser extensions pose significant privacy risks due to their access to sensitive user data, such as browsing history, local storage, payment information, and more [41]. These risks stem from the potential exploitation of extensions by malicious actors seeking unauthorized access to private information [13, 33]. Researchers have examined various aspects of these privacy risks such as detecting spying extensions [3], identifying inconsistencies in privacy practices [8], and finding privacy leaks [50]. There has also been some work to approach privacy issues from security perspective that include detecting vulnerable and malicious extensions [43], cloud-based security analysis [10], and experimental security analysis of sensitive data access [31, 5]. Chen et al. [9] revealed that 2.13% of Chrome browser extensions, representing over 60M users, can leak sensitive information such as browsing history, open tabs, passwords, and location data. Xie et al. [51] further investigated this issue, identifying hundreds of extensions that automatically extract user content from web pages [27]. Notable examples include popular extensions such as Paypal’s Honey, Capital One Shopping, Hola VPN, and Avira Safe Shopping, which exfiltrate browsing URLs and other sensitive information. Alarmingly, approximately 40% of Chrome extensions, including widely used ones, exhibit security vulnerabilities that could compromise private user data [33]. Moreover, Ling et al. [25] discovered that 92% of browser extensions engage in data collection practices that contradict their own privacy policies or stated practices. Many extensions were found to collect excessive data beyond what is necessary for their intended functionality. For instance, while the Chrome extension ‘InserLearning’ claimed to only collect a user’s name, Google account (email address), and Google profile image, it also gathered website content. Additionally, Carnus et al. [20] demonstrated that some extensions leak user data, such as email addresses, names, and phone numbers, to their developers and third-party entities.
@@
Privacy Issues in GenAI models. Past work in this space has mostly focused on discovering privacy issues resulting from the training phase of GenAI models. Researchers have demonstrated that LLMs can overfit to sensitive information within training datasets, potentially resulting in the unintended exposure of private data during interactions through APIs, browser extensions, or other interfaces [17, 23, 24, 21, 49, 37]. Such risks are amplified by the susceptibility of LLMs to various attack vectors, including model inversion attacks (the reconstruction of sensitive data from model outputs), adversarial attacks such as prompt injection, and data poisoning attacks involving the introduction of malicious data into training sets [53, 6, 55]. To mitigate these risks, prior research has investigated privacy-preserving methods such as differential privacy, federated learning, and secure multi-party computation aimed at reducing the exposure of sensitive user information [52]. Despite these advancements, critical gaps remain to prevent LLM models from learning sensitive attributes during the interaction phase. Staab et al. highlight that even limited data from users’ online activities on platforms like Reddit or Twitter enables an LLM to predict private attributes with 85% top-1 accuracy [45].
Figure 1: Our Threat model
@@
Comparison with prior works. Past research has primarily sought to study privacy-compromising data collection by extensions and analyze associated security vulnerabilities to protect sensitive user information from malicious actors. However, GenAI browser extensions integrated with modern LLMs introduce new risks. These include the potential exposure of user inputs, interactions, and browsing behavior to LLMs, which are highly efficient at inferring sensitive user attributes with high precision from seemingly benign interactions[45]. Our research builds on the findings of Staab et al. [45] by examining the compounded privacy risks of combining LLMs with browser extension capabilities, an area that has been largely overlooked. Our work is the first one focused at bridging this gap by performing a systematic audit of GenAI browser assistants. Additionally, past efforts have focused on analyzing leakage and sharing of data from user’s browser, without understanding how that data is used. GenAI browser assistants provide a unique vantage point to not just look at data leakage, but to also understand profiling and personalization capabilities based on the collected web-scale data. We investigate whether continuous tracking of user activities within their browser can help models associate memory with a user’s personal and behavioral data by periodically updating the context to personalize responses. Being a fairly new and developing space, it is important and timely to look at such unique risks posed by these assistants. We design a novel experimental framework to audit tracking, profiling, and personalization under 3 major functionalities – search, browse, and summarize – offered by 10 most popular search-based GenAI browser assistants.

### v2

@@ Abstract
Browser assistants have started to integrate powerful capabilities of GenAI in web browsers to offer functionalities such as question answering, content summarization, and agentic web navigation. These assistants, available today as browser extensions, raise significant privacy concerns because they can track detailed browsing activity (e.g., searches, clicks) and autonomously perform tasks such as form filling. In this paper, we analyze the design and behavior of GenAI browser extensions, focusing on how they collect, process, and share user data, and whether they profile users based on explicit or inferred demographic attributes and interests. We develop a novel prompting framework and perform network traffic analysis to audit the nine GenAI browser assistants for tracking, profiling, and personalization.
We find that GenAI browser assistants typically rely on server-side APIs rather than local models, and can be invoked automatically without explicit user interaction. GenAI browser assistants often collect and share full webpage content, including the HTML DOM and user form inputs in some cases, with their first-party servers. Some also share identifiers and user prompts with third-party trackers such as Google Analytics. This data collection and sharing happens even on pages containing sensitive information, such as health records or personal information such as names or social security numbers entered in a web form. Moreover, several GenAI browser assistants infer attributes (e.g., age, gender, income, interests) and use them to personalize responses across browsing contexts. Our findings show that GenAI browser assistants collect and share personal and sensitive information for profiling and personalization, highlighting the need for safeguards as they increasingly mediate web browsing.
@@ 1 Introduction
With recent advances in natural language processing, Large Language Models (LLMs) trained on massive datasets and built with billions of parameters have significantly improved their ability to understand context and generate content across modalities, including text, images, and videos [26]. As the foundational technology behind many Generative Artificial Intelligence (GenAI) systems, LLMs now power a wide range of applications such as chatbots and workflow automations. Popular search engines have also begun integrating LLMs. For instance, Google Search now displays Gemini-generated overviews in response to user queries [39], while Microsoft Bing’s Copilot uses models from OpenAI [25]. However, these search engines are limited to analyzing activity on their own platforms, restricting their visibility of a user’s browsing activity across the web. To address this limitation, GenAI browser assistants have emerged to use LLMs to enhance a user’s overall browsing experience across the web.
GenAI browser assistants are implemented as browser extensions, giving them access to nearly everything a user does in their browser. By operating as a browser extension, these assistants can track a user’s browsing activity across websites such as web pages visited, buttons clicked, search terms and other personal information typed into form fields. This browsing activity provides context that assistants can use to build detailed user profiles. While this enables personalized responses by these assistants, it also raises serious privacy concerns. For example, browsing data and user profiles can be collected, shared, and repurposed for other usecases such as online behavioral advertising. Despite high operational costs [19], most LLM platforms do not yet rely on ads for monetization. However, this shift is imminent [42, 47, 43].
It is important to understand the privacy risks posed by tracking, profiling, and personalization by GenAI-based browser assistants. Millions of users already use these assistants, yet little is known about how they are designed, what information they collect and share, and how they process it. To this end, we propose a novel framework to systematically audit GenAI browser assistants, guided by the following research questions:
@@
RQ1. How is the architecture of GenAI browser assistants designed? We analyze network traffic of these browser assistants to examine key design choices: (1) backend model capabilities, (2) response architecture, (3) context isolation across interactions, and (4) response variability.
RQ2. Do GenAI browser assistants collect and share user information? We analyze network traffic of these browser assistants to examine how user data is handled, focusing on: (1) implicit and explicit collection and sharing, (2) of data with first-party and third-party servers, and (3) differences across public and private online spaces.
RQ3. Do GenAI browser assistants profile users based on their browsing behavior to personalize responses? We develop a prompting framework to evaluate profiling and personalization, focusing on: (1) whether assistants infer user attributes such as location, age, gender, income, and interests from browsing activity, and (2) whether they use this information to personalize responses within and across browsing contexts.
We summarize our key contributions and findings below:
@@
$\bullet$ Architecture. We analyze nine popular GenAI browser assistants and find that all but one rely on server-side response generation; only a single assistant performs response generation on the client side. Some assistants are automatically invoked when a user enters a query on a search engine. Furthermore, seven assistants isolate context across browsing sessions and tabs, while the remaining two do not do so.
@@
$\bullet$ Tracking. Using assistants in private online spaces resulted in data collection ranging from partial webpage content to full DOM snapshots. Merlin was observed to extract form inputs, leading to the collection and sharing of sensitive information, including medical records from a university health portal (hem.ucdavis.edu), academic records from Canvas (from canvas.edu), and a Social Security number entered on an IRS website irs.gov). Sider and Merlin shared chat and user identifiers with google-analytics.com while TinaMind shared them with analytics.google.com. Merlin also shared raw user queries with Google Analytics, which could potentially be used for tracking and personalization across Google platforms. Chat histories were shared with first-party servers by four assistants. Moreover, Harpa and Copilot stored the complete chat history in the extension’s background service worker using IndexedDB, indicating that these histories persist across browsing sessions.
@@
$\bullet$ Profiling and Personalization. Two extensions (Monica, and Sider) demonstrated profiling based on all five user attributes (location, age, gender, income, and interests) and responded to both in-context and out-of-context personalization prompts. In contrast, Perplexity and TinaMind showed no strong evidence of profiling or personalization, while Harpa exhibited only in-context personalization behavior.
@@ 2 Background
We provide background of GenAI in Section 2.1 and its application in browser assistants in Section 2.2.
2.1 Generative AI
@@
How do GenAI systems work? Generative AI (GenAI) systems typically rely on large language models (LLMs), which operate by iteratively predicting the next token in a sequence based on the context of preceding tokens [52]. These models are built on transformer architectures, where a self-attention mechanism enables the model to assign varying importance to different parts of the input sequence [44]. When a user inputs a query, the model encodes the input into an embedding space and generates a probability distribution over possible next tokens. It then selects a token based on this distribution and repeats the process until the output is complete. This approach enables GenAI systems to provide coherent and context-aware responses to user input.
@@
Limitations and challenges of GenAI systems. GenAI systems face several challenges that affect their performance, usability, and broader applicability. Understanding these limitations is essential for evaluating how these systems function in practice. The first fundamental issue is their probabilistic nature. As described earlier, GenAI systems generate the next token based on a probability distribution, which can lead to varying outputs even for the same input. This variability is influenced by the temperature parameter, which controls the randomness of the output. Higher temperature values encourage the selection of less probable tokens, often producing more creative but potentially less coherent responses, while lower values lead to more deterministic and focused responses [26]. The second challenge is the resource intensive nature of training these models [6]. This means that these models cannot incorporate up-to-date information that was not part of the training process without access to live search capabilities. Finally, while general-purpose models are versatile, adapting them to domain-specific use cases, such as personalized assistants is non-trivial. These typically have access to only the user-submitted queries within the platform, which limits their context to what is explicitly included in the user prompt and makes personalization challenging.
@@ 2.2 Generative AI Browser Assistants
How do GenAI browser assistants function? Gen AI browser assistants aim to leverage capabilities of GenAI models to personalize user experience on the web. Implemented as browser extensions, these assistants function as wrappers around GenAI models such as OpenAI’s ChatGPT [36], Google’s Gemini [15], Meta’s Llama [23], etc. They monitor user activity on the web and use this context to query backend GenAI APIs and return personalized responses. Unlike foreground web applications, these assistants operate through background service workers that run continuously while the extension is active. They can inject content scripts into every page that a user visits. These content scripts may load additional libraries (e.g., jQuery) to facilitate actions like DOM manipulation, and can communicate with the background script to log and share user activity. This browser extension based architecture effectively grants the assistants broad access to everything a user does in the browser. By monitoring a user’s browsing activity across websites, browser assistants can offer in-the-moment suggestions tailored to specific browsing contexts. In summary, these assistants address many of the limitations of general-purpose GenAI systems and offer a more personalized, context-aware experience.
@@
What functionalities do GenAI browser assistants offer? GenAI browser assistants offer a range of features that we group into three broad categories: search-based assistants, content-based assistants, and automation assistants. Search-based assistants offer features such as natural language responses to search queries, support web page summarization, interactive Q&A with web pages, and text highlighting. Content-based assistants assist with content-specific tasks such as writing tasks, social media post generation, meeting transcript summarization, YouTube video explanation, SEO generation, etc. Automation assistants focus on streamlining workflows by supporting tasks such as auto-filling forms, coding and debugging agents, task scheduling, automated data extraction and scraping, voice-enabled assistance, etc.
2.3 Traditional vs. GenAI Browser Extensions
While browser extensions have long been supported in browsers, the integration of GenAI in browser extensions introduces new risks through expanded data collection and inference capabilities. To understand these risks, we highlight three key differences between traditional and GenAI-powered browser extensions:
Third-party model integration. GenAI extensions rely on third-party GenAI models to analyze a user’s browsing activity. This allows them to build rich user profiles by combining observed browsing behavior with inferred characteristics such as interests and intent. Distinct from traditional extensions, GenAI extensions can personalize responses based not only on what users explicitly do, but also on what the LLM infers about them through reasoning.
Extensive data collection. Traditional extensions typically collect limited, task-specific data such as URLs or selected DOM elements. For instance, a price comparison extension may only extract product names and prices from searches on a specific site like Amazon. On the contrary, GenAI extensions often collect and transmit full-page content or the entire DOM.
Memory. Many GenAI browser assistants include a memory component that persists across navigations, sessions, or tabs. This enables longitudinal tracking and cross-context profiling, which is not common in traditional extensions.
3 Related Work
@@
In this section, we review related literature along two dimensions: privacy issues in browser extensions and privacy issues in GenAI. We conclude by discussing how our work compares to and builds upon prior research in this space.
@@
Privacy issues in browser extensions. While browser extensions add to functionality, they pose significant privacy risks due to their access to user data such as browsing history, local storage, payment information, and more [38]. These risks stem from the potential exploitation of extensions by malicious actors seeking unauthorized access to user data [12, 31]. Prior research has examined various aspects of these privacy risks such as detecting spying extensions [1], identifying inconsistencies in privacy practices [7], and finding data leaks [48]. There has also been some work from the security perspective, focusing on detecting vulnerable and malicious extensions [40], enabling cloud-based security analysis [10], and security analysis of sensitive data access [29, 4]. Chen et al. [8] found that 2.13% of Chrome browser extensions, representing over 60 million users, leak sensitive information such as browsing history, open tabs, passwords, and location data. Xie et al. [49] further further identified hundreds of extensions that automatically extract user content from web pages [24]. Well-known examples include popular extensions like PayPal’s Honey, Capital One Shopping, Hola VPN, and Avira Safe Shopping, which exfiltrate URLs and other user data. Alarmingly, around 40% of Chrome extensions, including the widely used ones, exhibit security vulnerabilities that could compromise user data [31]. Moreover, Ling et al. [22] found that 92% of browser extensions collect data in ways that contradict their stated privacy policies. Many extensions were found to collect excessive data beyond what was necessary for their intended functionality. For instance, while the Chrome extension InserLearning claimed to only collect a user’s name, Google account email, and profile image, but was also found to collect full webpage content. Additionally, Carnus et al. [17] demonstrated that some extensions leak user data, such as email addresses, names, and phone numbers, to their developers and third parties.
@@
Privacy issues in GenAI models. Past work in this space has mostly focused on investigating privacy issues resulting from the training phase of GenAI models. Researchers have demonstrated that LLMs can overfit to sensitive information within training datasets, potentially resulting in the unintended exposure of private data during interactions through APIs, browser extensions, or other interfaces [16, 20, 21, 18, 46, 37]. Such risks are amplified by the susceptibility of LLMs to various attack vectors, including model inversion attacks (reconstruction of sensitive data from model outputs), adversarial attacks (e.g., prompt injection), and data poisoning attacks (introduction of malicious data into training sets) [51, 5, 53]. To mitigate these risks, prior research has investigated privacy-preserving methods such as differential privacy, federated learning, and secure multi-party computation aimed at reducing the exposure of sensitive user information [50]. There remain critical gaps to prevent GenAI models from learning sensitive attributes during the inference phase. Staab et al. show that even limited data from users’ online activities on platforms like Reddit or Twitter enables an LLM to infer a user’s personal attributes with 85% accuracy [41].
@@
Comparison with prior work. Past research has primarily focused on privacy-invasive data collection by browser extensions and the associated security vulnerabilities that expose user data to malicious actors. However, GenAI browser extensions integrated with LLMs introduce new risks, as discussed in Section 2.3. These risks include potential exposure of browsing activity, form inputs and other interaction to LLMs, which are highly effective at inferring sensitive user attributes from seemingly benign interactions [41]. Our work builds on the findings of Staab et al. [41] by examining the compounded privacy risks of combining LLMs with browser extension capabilities, an area that has been understudied. To the best of our knowledge, ours is the first systematic audit focused specifically on GenAI browser assistants. While prior work has analyze collection and sharing of user data, they have not investigated how that data is subsequently used. GenAI browser assistants are a new target for privacy analysis, allowing us to investigate not just data collection and sharing but also how it may be used for profiling and personalization. We investigate whether continuous tracking of user activities within the browser allows assistants to associate memory with a user’s personal and behavioral traits, using updated context to personalize responses over time. Being a fairly new and developing space, it is both important and timely to investigate the unique privacy risks they pose. To that end, we design a novel experimental framework to audit tracking, profiling, and personalization across three major functionalities (search, browsing, and summarization) offered by nine of the most widely used GenAI browser assistants.